OpenStudy (anonymous):

all real values of x find them in 2/x+1=x-2/2? and if its to much to ask how do u work this out?

OpenStudy (lalaly):

cross multiply 2*2 = (x+1)(x-2) 4=x^2 -2x +x -2 x^2-x -6 =0 (x-3)(x+2)=0 x=3 x=-2

OpenStudy (anonymous):

(x-3)(x+2) = 0 x=3, x=-2

OpenStudy (anonymous):

You can multiply up the bottom of each side so \[2(2)=(x-2)(x+1)\] which multiplied out will give you \[4=x ^{2}-x-2\] Rearranged \[x ^{2}-x-6=0\] Solved for x x=3 x=-2

OpenStudy (anonymous):

Thank u so much omg ive been at it for about 30 min omg thanks